# Bidirectional predictive coding

### v1

@@ 3.2 Flexible learning
bPC can be trained both in supervised and unsupervised settings. In all cases, neurons in intermediate layers (layer $2$ to $L-1$) are un-clamped and evolve according to the neural dynamics given in equation 5. In supervised settings, the first layer $x_{1}$ is clamped to the input, while the top layer $x_{L}$ is clamped to the target labels. In unsupervised settings, $x_{L}$ is left un-clamped, allowing bPC to learn compressed representations of the input. A mixed setting is also possible, where only a subset of neurons in $x_{L}$ are clamped to label information, while others remain free. In this setting, the model can jointly infer labels and learn an associated compressed representation.
@@ 4 Experiments
4.1 Supervised classification and generation
@@
Figure 3: bPC accurately classifies and generates class average images on MNIST and Fashion-MNIST. A: Training set-up. The models are trained with $x_{1}$ fixed to images and $x_{L}$ fixed to labels. B: Classification accuracy of models. C: Examples of the generated images conditional on class labels (left) and RMSEs between generated images and mean images of each class (right).
Firstly, we demonstrate that a single bPC model simultaneously performs discriminative and generative tasks. In this experiment, bPC learns to associate images and labels, such that it can classify an input image and generate an average image of a class when a label is provided. We compare bPC with discPC, genPC, hybridPC, and their backpropagation equivalents on MNIST and Fashion-MNIST datasets for the same architecture with two hidden layers (256 neurons each). Additional PC models are evaluated in the supplementary material, SM. During training, we clamped layer $x_{1}$ to images and layer $x_{L}=x_{4}$ to their corresponding one-hot labels (Figure 3A). Post-training, we evaluated discriminative performance by measuring classification accuracy given only the image. Specifically, we fixed layer $x_{1}$ to the image and inferred labels at layer $x_{L}$ after 100 inference steps. Generative performance was assessed by the root mean squared error (RMSE) between the image generated from each class label and the corresponding average image for that class. We obtained generated images in $x_{1}$ after fixing $x_{L}$ to a class label and performing a 100-step iterative inference. A detailed training and evaluation procedure for all experiments, including hyperparameter selection, is provided in SM.
@@
A comparison of classification accuracy and generative RMSE across models is shown in Figure 3B. bPC achieves a classification accuracy comparable to discPC and discBP for both datasets. Conversely, genBP and hybridPC show lower accuracy. Regarding generative performance, bPC exhibits RMSE scores similar to genPC, hybridPC, and genBP, while discPC demonstrates significantly higher generation errors. Generated images from discPC contain minimal class-relevant features, whereas bPC and other models produce clear, representative images for each class (Figure 3C).
These results confirm previously reported results that unidirectional PC models excel only in their specialized domain: discPC fails in generation due to non-unique learning minima [9], while genPC and hybridPC have a poor classification accuracy [19]. In contrast, bPC performs well in generative and discriminative tasks simultaneously. Notably, bPC uses only half the neurons needed to maintain two pathways for generative and discriminative tasks separately, which can be more energy efficient.
We also found that discriminative weighting ($\alpha_{disc}$) needs to be set higher than the generative weighting due to the larger magnitude of top-down prediction errors. The exploration into the impact of $\alpha_{disc}/\alpha_{gen}$ is presented in SM.
4.2 Unsupervised representation learning
Figure 4: bPC outperforms other PC models at learning low-dimensional representations. A: Training set-up, where only $x_{1}$ is clamped to input images. B: Reconstruction RMSE across models and datasets. C: Example reconstructions.
@@ 4.2 Unsupervised representation learning
Next, we show that bPC learns compressed representations in the absence of supervision. We evaluated the representation learning performance of bPC, genPC, hybridPC, and their BP equivalents across tasks of increasing complexity, using MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. discPC was not evaluated due to its inability to perform unsupervised learning. We trained models where only layer $x_{1}$ was clamped to input images, leaving all other layers free (see Figure 4A). MNIST and Fashion-MNIST models had two hidden layers (256 neurons each) and a representation layer $x_{L}$ of 30 neurons. Training was constrained to 8 inference steps between weight updates to assess the models’ capacity for fast inference. CIFAR models had five convolutional layers and a representation layer of 256 neurons, trained with 32 inference steps. Activity decay was used at layer $x_{L}$ to stabilize learning and constrain representations. After training, representations were obtained in $x_{L}$ after 100 inference steps for a given input image in $x_{1}$. Images were reconstructed using 100 inference steps after clamping $x_{L}$ to a representation and reinitialising layers $x_{1}$ to $x_{L-1}$. We measured the RMSE between the original images and their reconstructions from representations.
Figure 4B shows that bPC consistently outperforms genPC by achieving a lower RMSE. On MNIST and Fashion-MNIST, bPC matches the performance of hybridPC. On more complex datasets such as CIFAR-10 and CIFAR-100, bPC significantly surpasses hybridPC. In these cases, it approaches the performance level of BP-based models. These results are also illustrated in Figure 4C with example image reconstructions for each model. The superior representation learning capability of bPC over genPC suggests that a feedforward initialization through the bottom-up pathway in bPC serves as a fast amortised inference that nudges its initial neural activities towards the optimal state. This is particularly useful when inference steps are limited. The advantage of bPC over hybridPC is likely due to the active involvement of the bottom-up weights $V_{l}$ in the inference dynamics, which delivers sensory information to the latent neurons during iterative inference. This becomes more important when the sensory inputs are more complex, such as in the CIFAR datasets. The local nature of computations in bPC, which leads to lower learning efficiency but more plausible implementation, could underlie its slightly worse performance than BP-based models.
4.3 Combined supervised and unsupervised learning
@@
Here, we integrate the supervised and unsupervised settings described previously to show that bPC can simultaneously develop discriminative capabilities and compressed representations. We trained bPC, hybridPC and their BP-based equivalents on the MNIST, Fashion-MNIST, and CIFAR-10 datasets. During training, we clamped layer $x_{1}$ to input images and partially clamped the top-layer $x_{L}$ to one-hot labels, while leaving the remaining neurons of $x_{L}$ free to learn complementary representations (see Figure 5A). Consistent with previous sections, MNIST and Fashion-MNIST models comprised two hidden layers of 256 neurons each, with layer $x_{L}$ composed of 40 neurons—10 dedicated to labels and 30 for representations. The CIFAR-10 model employed a convolutional architecture, with four hidden convolutional layers with max-pooling and a latent layer $x_{L}$ containing 266 neurons (10 for labels, 256 for representations). Additionally, activity decay was applied to the representational neurons in $x_{L}$. Classification accuracy was measured by providing only images (as in Section 4.1), while generative quality was assessed by reconstructing images from their representations (as in Section 4.2). The key difference is that images are reconstructed by clamping $x_{L}$ to the inferred representation and the image label.
Figure 5B reveals that bPC achieves classification accuracy on par with BP models across all datasets. Furthermore, it significantly surpasses hybridPC, particularly on CIFAR-10, where hybridPC exhibits more than 45% lower accuracy than bPC. In terms of generative quality, Figure 5D shows that bPC achieves reconstruction errors similar to hybridPC across all datasets. The reconstructed examples in Figure 5C illustrate that bPC goes beyond generating class-average images, capturing features such as shape, color, and spatial locations (e.g., it generates ‘4’s in different styles). However, fine details are absent in bPC’s CIFAR-10 reconstructions due to artefacts introduced by max-pooling operations in the discriminative bottom-up pathway, affecting the generative top-down reconstruction process. Consequently, bPC shows stronger grid-like artefacts than hybridPC, which uses only max-pooling for initialization, and substantially more than BP models with separate generative and discriminative layers. While removing max-pooling removes artefacts, it reduces classification accuracy (see SM). Nonetheless, even without max-pooling, bPC’s classification accuracy significantly outperforms hybridPC, while generative performance remains comparable. Future research could refine bPC’s architecture to balance classification accuracy with artefact-free generation.
@@
Figure 5: bPC is the only PC model that can jointly learn low-dimensional representations of images and accurately classify them. A: Training set-up, where the latent layer is only partially clamped to class labels. B: Classification accuracy. C: Example reconstructions on MNIST and CIFAR10. D: Reconstruction RMSEs.
@@
4.4 Shared latent layers prevent biased or overconfident models
To understand why bPC is able to simultaneously perform both discriminative and generative tasks well, we examine its energy landscape (Eq 4). We first consider the XOR task, a simple, $2$-dimensional nonlinear classification problem where the energy landscape can be visualised. We compare bPC’s energy landscape to those developed by individual genPC and discPC, to investigate the advantage of shared, rather than independent, latent layers. We trained bPC (Figure 6A bottom), genPC and discPC (Figure 6A top), each with two hidden layers of 16 neurons, by clamping $x_{1}$ to the $2$-dimensional inputs and $x_{4}$ to the scalar outputs. In this set-up, bPC has an equal number of parameters to that of genPC and discPC combined, but half the number of neurons. The resulting energy landscapes after training are shown in Figure 6B, where discPC exhibits broad regions of low energy, indicating overconfidence across a wide range of input configurations, including potentially OOD inputs. genPC, on the other hand, develops energy minima at the mean of each class, failing to capture the true structure of the XOR data. Notably, when discPC and genPC are combined into a single model with separate objectives, the resulting energy landscape retains both problems: broad and biased minima. In contrast, bPC learns sharp, well-localized minima around data points.
We then examined the generalisability of these observations, by evaluating the ability of the models considered in Section 4.1 to generate MNIST samples. Specifically, we clamped the topmost layer $x_{L}$ to a specific digit label, randomly initialized the input layer $x_{1}$, and performed inference updates until the model reached energy values equal or lower to the lowest 10% of equilibrium energies observed on MNIST test images. For a combined discPC-genPC model, we required the generated samples to satisfy this low-energy criterion for both components. This set-up serves as a sampling process, allowing us to inspect images that are considered as highly likely by the models. Figure 6C shows that samples generated by bPC are substantially more realistic than those from the combined discPC-genPC model, which often include excessive background noise and poorly defined digit shapes. To quantify sample quality, we used the Inception Score, that assesses how easily the generated digits can be labeled by a classifier [32], and the Fréchet Inception Distance (FID), which measures image quality and diversity [33]. bPC significantly outperformed the combined model, with higher Inception Scores (bPC: $6.05^{\pm 0.17}$ vs. combined: $3.62^{\pm 0.03}$, mean ± SEM over 5 seeds) and lower FID scores (bPC: $44.4^{\pm 2.2}$ vs. combined: $140.5^{\pm 2.1}$). We repeated the comparison with a BP-based combined model (discBP + genBP), and once again bPC produced higher-quality samples (see SM).
These results provide a detailed mechanistic explanation for the observations in Section 4.1. discPC performs poorly in image generation by confidently assigning low energy to nonsensical, OOD images with its broad energy minima. genPC recognizes only the mean image of each class as low-energy, thus relying on a template-matching mechanism where each new pattern is classified based on its distance to the mean class images. bPC, on the other hand, develops a more selective energy landscape as these two energy functions ($E_{disc}$ and $E_{gen}$) are jointly minimised and act as mutual regularisations on each other. This can be observed from bottom panel of Figure 6B, where $E_{gen}$ in bPC is ‘stretched’ by $E_{disc}$ to also include the individual data points, whereas $E_{disc}$ is ‘squeezed’ by $E_{gen}$ to have smaller, more dedicated minima.
@@ 4.5 Learning in biologically relevant tasks
Figure 7: bPC performance in multimodal learning and robustness against missing information. A: Bimodal bPC and genPC. Note that bimodal bPC is fully equivalent to unimodal bPC in Section 4.1; it is bent here to serve visual purposes. B: Classification accuracy of bimodal PC models. C & D: Generation RMSE and examples. E: Model evaluation with partially occluded sensory inputs. F & G: Classification accuracy against percentage of missing pixels with MNIST and Fashion-MNIST respectively. H: Activity of $x_{1}$ after inference for genPC and bPC.
First, we examine bPC in multimodal learning, which is a typical way the brain forms associations [34]. When simulating supervised learning tasks, the information on the label is typically provided to a top layer for simplicity, however, when humans acquire information, the category is often represented in another modality. For example, when a child learns names of animals, these names are spoken by a parent and processed via the auditory cortex (Figure 7A). To assess multimodal capabilities, we compared bPC and genPC on a bimodal task using the MNIST and Fashion-MNIST datasets. We trained hierarchical models with one latent layer connected to two separate input layers (Figure 7A), one clamped to a one-hot vector of label and the other to the corresponding input image. The bPC model included bottom-up and top-down predictions between each input layer and the latent layer, whereas genPC incorporated only top-down predictions from the latent to the input layers.
After training, we evaluated cross-modal information transfer by measuring classification accuracy and average image reconstruction error. We provided input to one modality and ran iterative inference. Next, we compared the activity of the modality without input against either the correct label or the class-averaged image. Figure 7B, C and D illustrates that the bimodal bPC model significantly outperforms genPC in both classification accuracy and image reconstruction performance. This result is expected as bimodal bPC is equivalent to the bPC model introduced in Section 4.1 due to the bottom-up and top-down symmetry, which has been proven successful in both tasks. In other words, hidden layers of bPC naturally model associative cortical areas. On the other hand, genPC has to be purposely re-structured to model the multimodal association cortex, where one pathway predicts the image (similar to genPC) and the other predicts the label (similar to discPC). Consequently, bimodal genPC inherits limitations from both model types, resulting in poor classification and generation.
Next, we test the robustness of bPC to partially missing information, as the brain is known to retain important cognitive functions such as categorisation, even with incomplete or occluded visual scenes [35]. We used trained models (bPC, genPC, discPC, and discBP) from section 4.1 to assess their classification accuracy under increasing levels of missing pixels (from 0% to 90%) on MNIST and Fashion-MNIST. We fix the activity of observed neurons in $x_{1}$ to observed pixels and leave unobserved neurons free, initialized to zero activity as shown in figure 7E. To accommodate the slower inference convergence with missing data, we allow extended inference durations before evaluation.
As shown in Figure 7F and G, bPC maintains high classification accuracy even with up to 80% of pixels missing. genPC also retains reasonable performance, though its overall accuracy is lower than bPC across levels of occlusion. By contrast, discPC and discBP exhibit sharp performance drops beyond 50% missing input. bPC’s robustness stems from its top-down predictions: they actively fill in missing input through inference, reconstructing inputs by integrating top-down generative priors and bottom-up evidence. Figure 7H shows example input images where missing inputs are filled in by bPC. genPC shares this capacity but suffers from weak discriminative power. In contrast, discPC and discBP cannot infer missing information because they lack the top-down generative mechanism.

### v2

@@ 3.2 Flexible learning
bPC can be trained both in supervised and unsupervised settings. In all cases, neurons in intermediate layers (layer $2$ to L1L-1) are un-clamped and evolve according to the neural dynamics in equation 4. In supervised settings, the first layer $x_{1}$ is clamped to the input, while the top layer $x_{L}$ is clamped to the target labels. In unsupervised settings, $x_{L}$ is left unclamped, allowing bPC to learn compressed representations of the input. A mixed setting is also possible, where only a subset of neurons in $x_{L}$ are clamped to label information, while others remain free. In this setting, the model can jointly infer labels and learn an compressed representation.
@@ 4 Experiments
4.1 bPC performs simultaneously well in classification and generation
@@
Figure 3: bPC accurately classifies and generates class average images on MNIST and Fashion-MNIST. A: Training set-up. The models are trained with $x_{1}$ fixed to images and $x_{L}$ fixed to labels. B: Classification accuracy of models. C: Examples of the generated images conditional on class labels (left) and RMSEs between generated images and mean images of each class (right). Error bars denote the standard error of the mean (s.e.m.) across 5 seeds.
In this experiment, we assessed bPC’s capacity for both classification (discriminative) and class-conditional image generation (generative). We compared bPC with discPC, genPC, hybridPC, and their backpropagation equivalents on MNIST and Fashion-MNIST, using identical architectures with two hidden layers of 256 neurons each. Additional baselines, including bPC with shared bottom-up and top-down weight as proposed by Qui et al (2023), are provided in the supplementary material (SM) C.
During training, the input layer $x_{1}$ was clamped to images and the output layer $x_{L}=x_{4}$ to their corresponding one-hot labels (Figure 3A). After training, discriminative performance was assessed by fixing $x_{1}$ to an input image and inferring the label at $x_{L}$ over 100 inference steps. Generative performance was evaluated by clamping $x_{L}$ to a class label and measuring the root mean squared error (RMSE) between the inferred image at $x_{1}$ (after 100 inference steps) and the average image for that class.
@@
The results are shown in Figure 3B. On both datasets, bPC achieved classification accuracy comparable to discPC and discBP, while genBP and hybridPC performed less well. For generation, bPC obtained RMSE scores similar to genPC, hybridPC, and genBP, whereas discPC exhibited much higher errors. Visualizations in Figure 3C further illustrate these differences: discPC generated images with little class-relevant structure, while bPC and other bidirectional models produced clear and representative samples.
These findings align with previous reports that unidirectional PC models excel only in their specialized domain (Sun & Orchard, 2020; Tscshantz et al., 2023), and demonstrate that bPC can integrate both discriminative and generative capabilities within a single framework. Notably, this integration uses the same number of error neurons but only half the value neurons required for maintaining separate unidirectional pathways for the two tasks. This indicates that bPC is more energy-efficient, especially as error signals can be handled in the dendrites of neurons rather than by a separate neuron population (Mikulasch et al., 2023). We observed that the discriminative weighting parameter ($\alpha_{disc}$) must be set higher than the generative weighting ($\alpha_{gen}$) due to the larger magnitude of top-down prediction errors. An exploration of the trade-off between $\alpha_{disc}$ and $\alpha_{gen}$ is provided in SM D.
4.2 bPC performs unsupervised representation and distribution learning
Figure 4: bPC matches the performance of hybridPC at unsupervised learning. A: Training set-up, where only $x_{1}$ is clamped to input images. B: Reconstruction RMSE from representations. C: Example reconstructions. D: Linear decoding accuracy from representations. E: Fréchet Inception Distance of samples generated by models with stochastic dynamics, with examples shown. T denotes the number of inference steps used during training. Error bars show s.e.m. across 5 seeds.
@@
Next, we show that bPC learns compressed representations and data distributions in the absence of supervision. We compared bPC, genPC, hybridPC, and their backpropagation (BP) equivalents on MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100, excluding discPC since it cannot perform unsupervised learning. We clamped only $x_{1}$ to input images, leaving all other layers free (see Figure 4A). For MNIST and Fashion-MNIST, models used two hidden layers (256 neurons each) and a 30-neuron representation layer $x_{L}$, trained with only 8 inference steps per update to test fast inference. CIFAR models had five convolutional layers and a 256-neuron representation layer, trained with 32 inference steps. An activity decay term at $x_{L}$ stabilized learning and regularized representations. After training, representations were obtained from $x_{L}$ using 100 inference steps. We evaluated these representations in two ways. First, reconstruction quality was measured by reinitializing layers $x_{1}$-$x_{L-1}$, clamping $x_{L}$, and running 100 inference steps, with RMSE computed against the input. Second, linear readout classification accuracy was measured from $x_{L}$.
To assess distribution learning, we compared bPC, genPC, hybridPC, and a VAE (Kingma et al., 2013) on MNIST, extending all PC models with stochastic dynamics to infer full posteriors ($\sigma^{2}=1$ in equation 4, see SM A.5 for more details). Architectures matched those above, trained with 50 or 250 inference steps. Models were evaluated using Fréchet Inception Distance (FID). Samples were generated ancestrally by following top-down predictions starting from sampling $x_{L}$ from its Gaussian prior and propagating downward through conditional Gaussians.
As shown in Figure 4B-E, bPC consistently outperformed genPC, yielding lower reconstruction RMSE, higher decoding accuracy, and better FID. It matched hybridPC across most tasks, and on CIFAR datasets significantly surpassed it in reconstruction error, approaching BP-based baselines. Reconstructions and generated samples (Figures 4C, E) illustrate these differences.
These results highlight that bPC outperforms genPC when inference steps are limited, and matches hybridPC despite the latter’s amortized pathway. This suggests that bPC’s bottom-up pathway also serves as amortized inference, rapidly initializing neural activities toward the optimal state. The slight edge of bPC over hybridPC likely arises from the active involvement of bottom-up weights $V_{l}$ during iterative inference, which continuously deliver sensory information to latent neurons. This effect becomes more pronounced with complex inputs such as CIFAR.
4.3 bPC performs combined supervised and unsupervised learning
In this experiment, we combined the supervised and unsupervised settings described above to test whether bPC can simultaneously develop discriminative capabilities and compact representations. This setting is motivated both computationally and biologically: real-world learning rarely occurs in isolation, and cortical circuits appear to integrate categorical supervision (e.g., from higher-order areas) with unsupervised structure learning from sensory input.
@@
We trained bPC, hybridPC, and their BP equivalents on MNIST, Fashion-MNIST, and CIFAR-10. During training, the input layer $x_{1}$ was clamped to images, while the top layer $x_{L}$ was partially clamped to one-hot labels, leaving the remaining neurons free to learn complementary representations (Figure 5A). For MNIST and Fashion-MNIST, models had two hidden layers of 256 neurons each, and $x_{L}$ comprised 40 neurons (10 for labels, 30 for representations). For CIFAR-10, we used a convolutional architecture with four hidden convolutional layers and a 266-neuron latent layer ($x_{L}$, with 10 label and 256 representation neurons). Activity decay was applied to the representational subpopulation in $x_{L}$ to regularize learning. Classification accuracy was evaluated as in Section 4.1, by presenting only images and inferring labels at $x_{L}$. Generative quality was assessed as in Section 4.2, but with reconstructions conditioned jointly on the inferred representation and the label.
@@
Figure 5: bPC is the only PC model that can jointly learn low-dimensional representations of images and accurately classify them. A: Training set-up, where the latent layer is only partially clamped to class labels. B: Classification accuracy. C: Example reconstructions on MNIST and CIFAR10. D: Reconstruction RMSEs. Error bars show s.e.m. across 5 seeds.
Figure 5B reveals that bPC achieves classification accuracy on par with BP models across all datasets. Furthermore, it significantly surpasses hybridPC, particularly on CIFAR-10, where hybridPC exhibits more than 45% lower accuracy than bPC. In terms of generative quality, Figure 5D shows that bPC achieves reconstruction errors similar to hybridPC across all datasets. The reconstructed examples in Figure 5C illustrate that bPC goes beyond generating class-average images, capturing features such as shape, color, and spatial locations (e.g., it generates ‘4’s in different styles). However, fine details are absent in bPC’s CIFAR-10 reconstructions due to artefacts introduced by max-pooling operations in the discriminative bottom-up pathway, affecting the generative top-down reconstruction process. Consequently, bPC shows stronger grid-like artefacts than hybridPC, which uses only max-pooling for initialization, and than BP models with separate generative and discriminative layers. While removing max-pooling removes artefacts, it reduces classification accuracy (see SM E). Nonetheless, even without max-pooling, bPC’s classification accuracy significantly outperforms hybridPC, while generative performance remains comparable. Future research could refine bPC’s architecture to balance classification accuracy with artefact-free generation.
@@
4.4 bPC’s shared latent layers prevent biased or overconfident energy landscape
@@
In this experiment, we investigated why bPC can perform both discriminative and generative tasks effectively, focusing on the role of its shared latent layers in shaping the energy landscape (Eq. 3). As a toy example, we considered the XOR problem, where the landscape can be directly visualized. We trained bPC, discPC, and genPC (each with two hidden layers of 16 neurons) by clamping $x_{1}$ to 2D inputs of XOR and $x_{4}$ to scalar outputs (Figure 6A). bPC used the same number of parameters as discPC and genPC combined but half the neurons.
An ideal XOR landscape has well-localized minima only at valid input-label pairs ([0,1] and [1,0] for one class; [0,0] and [1,1] for the other). Figure 6B top shows that discPC instead develops broad low-energy regions, indicating overconfidence even for implausible or out-of-distribution (OOD) inputs. genPC collapses each class into a single mean, failing to capture the true structure. A combined discPC+genPC model inherits both flaws. In contrast, bPC learns sharp, class-specific minima centred on valid inputs (Figure 6B bottom).
To test generality, we evaluated the models from Section 4.1 by sampling MNIST digits. We clamped $x_{L}$ to a label, initialized $x_{1}$ randomly, and iterated inference until reaching energies within the lowest 10% of those observed on test images. For the combined discPC+genPC model, we required the generated samples to satisfy this low-energy criterion for both components. This set-up serves as a sampling process, allowing us to inspect images considered as highly likely by the models. As shown in Figure 6C, bPC produced realistic digits, while the combined model yielded poorly formed shapes. Quantitatively, bPC achieved higher Inception Scores ($6.05\pm 0.17$ vs. $3.62\pm 0.03$) and lower FID ($44.4\pm 2.2$ vs. $140.5\pm 2.1$). Similar trends held for BP-based baselines (see SM F).
These results clarify the observations in Section 4.1. discPC fails at generation because it produces broad, overconfident minima that admit many nonsensical and OOD inputs, leading to the noisy conditional generations in Figure 3 and in Sun & Orchard (2020). genPC, by contrast, learns narrow minima around class means, which rejects OOD inputs but identifies deviations from training-set mean as high-energy states, ultimately sacrificing discriminative precision. bPC integrates the two: bidirectional predictions effectively regularize each other, sharpening discriminative minima while anchoring them to specific data points through the generative process. Shared latent layers thus prevent solutions overly biased to training mean or those that are overconfident, yielding superior performance in both classification and generation.
@@ 4.5 Learning in biologically relevant tasks
Figure 7: bPC performance in a bimodal architecture and robustness against missing information. A: Bimodal bPC and genPC. Note that bimodal bPC is fully equivalent to unimodal bPC in Section 4.1; it is bent here to serve visual purposes. B: Classification accuracy of bimodal PC models. C & D: Generation RMSE and examples. E: Model evaluation with partially occluded sensory inputs. F & G: Classification accuracy against percentage of missing pixels with MNIST and Fashion-MNIST respectively. H: Activity of $x_{1}$ after inference for genPC and bPC. Error bars and shaded region show s.e.m. across 5 seeds.
In this section, we demonstrate that bPC outperforms other predictive coding models in two biologically relevant scenarios: (1) learning in a model architecture with two input streams analogous to two sensory modalities, and (2) classifying images from partially occluded inputs, similar to how human vision includes regions with missing sensory information, such as the retinal blind spot.
Bimodal model architecture.
The brain often develops neural representations through associations across modalities, such as linking spoken names to visual objects (Rosen et al., 2018). To test whether bPC can form such associations, we trained bPC and a bimodal variant of genPC on MNIST and Fashion-MNIST, with one latent layer connected to two inputs: an image and its one-hot label (Figure 7A; note that bPC is naturally bimodal and thus requires no restructuing). bPC incorporated both bottom-up and top-down connections for each modality, whereas genPC relied only on top-down pathways. After training, we evaluated cross-modal transfer by providing input to one modality and measuring inference quality in the other, classification accuracy for labels and RMSE for reconstructed images. As shown in Figure 7B-D, bPC significantly outperformed bimodal genPC on both tasks. This result is consistent with Sections 4.1: bPC’s bidirectional pathways naturally support associative coding, while genPC must be restructured to handle multimodal inputs, where one pathway predicts the image (similar to genPC) and the other predicts the label (similar to discPC) and inherits the weaknesses of both genPC and discPC.
Robustness to missing information.
The cortex can recognize objects even when sensory input is incomplete or occluded (Komatsu, 2006). We therefore tested the classification performance of the trained models from Section 4.1 (bPC, genPC, discPC, and discBP) under progressive occlusion of MNIST and Fashion-MNIST inputs, masking up to 90% of pixels (Figure 7E). Observed pixels were clamped in $x_{1}$, while missing ones were left free and initialized to zero. We then run inference on the missing pixels and in latent layers, with extended iterations to accommodate slower convergence. Results (Figure 7F-G) show that bPC maintained high classification accuracy even with 80% missing pixels. genPC also retained some robustness but at consistently lower accuracy. In contrast, discPC and discBP collapsed beyond 50% occlusion. Visualizations in Figure 7H reveal that bPC actively reconstructed missing inputs via top-down priors, integrating them with observed evidence. genPC shared this generative capacity but lacked strong discriminative accuracy, while purely discriminative models could not compensate for missing information at all.
4.6 Generative and discriminative processing in deep models
To investigate whether bPC scales effectively to deeper models and more complex datasets, we trained bPC, discPC and discBP using the following architecture-dataset combinations: VGG-5 on CIFAR-10, VGG-9 on CIFAR-100, and VGG-16 on Tiny-ImageNet. After training, we evaluated discriminative performance by measuring the model’s classification accuracy. Additionally, we assessed generative capabilities by evaluating classification accuracy when 30% and 50% of the input pixels were missing, following the methodology outlined in Section 4.5. To efficiently simulate the predictive coding models in this experiment, we employed error optimisation (Goemaere et al., 2025). This approach prevents energy decay in predictive coding models and enables the training of larger architectures. Refer to SM I for more implementation details.
Figure 8 confirms that bPC successfully combines discriminative and generative processing in deep networks. bPC achieves a classification accuracy comparable to discPC and discBP when full images are presented and significantly outperforms discPC and discBP on images with missing input information. For example, when 50% of pixels are missing, bPC classifies more than 60% more accurately than discPC and discBP on CIFAR-10 (Figure 8A). This improvement stems from bPC’s generative processing, which fills in missing information. Figure 8D illustrates this effect: after inference, the activity of the input layer $x_{1}$ reveals that neurons lacking inputs have been updated to predict the missing values. Once the missing information is inferred, bPC classifies images accurately. Overall, even in larger models, bPC effectively balances discriminative and generative performance by jointly minimising bottom-up and top-down prediction errors.
Figure 8: bPC’s discriminative and generative properties scale to deep networks. A & B & C: Classification accuracy against the percentage of missing pixels for bPC, discPC and discBP on the architecture-dataset combinations: VGG-5 on CIFAR-10, VGG-9 on CIFAR-100, and VGG-16 on Tiny-ImageNet. D: Activity of input neurons $x_{1}$ after inference for bPC trained on Tiny-ImageNet. Error bars show s.e.m. across 5 seeds.